OpenStudy (anonymous):

Find the standard form of the parabola with a focus of (-2, 0) and a vertex of (0, 0).

OpenStudy (anonymous):

y^2 = -8x

OpenStudy (anonymous):

you sure ?????

OpenStudy (anonymous):

yes miss

OpenStudy (amistre64):

...... maybe lol

OpenStudy (amistre64):

y^2 = 4ax ; and 'a' = distance from vert to foci = -2 y^2 = -8x .... got lucky that time lol

OpenStudy (amistre64):

focus is in the opposite direction and remains onthe x axis...right?

OpenStudy (amistre64):

ack.. its right .... i gotta read these things bette

OpenStudy (anonymous):

ok cooooollllllllll thanks :)

OpenStudy (amistre64):

im watching the owl; gonna take him down lol

OpenStudy (amistre64):

...hope you bought a calendar, cause your days are numbered .... maybe